OpenStudy (anonymous):

solve for x please!!!!!! -14x=-11x+3

OpenStudy (anonymous):

14x+11x=3 25x=3 x=3/25

OpenStudy (anonymous):

its( -14x) @Parvathysubhash

OpenStudy (anonymous):

( -14x)=-11x+3

OpenStudy (anonymous):

srry -14x+11x=3 3x=3 x=3/3=1

OpenStudy (anonymous):

so x=1

OpenStudy (anonymous):

you were wrong............and look at the equation.....but i got it wrong its not one

OpenStudy (anonymous):

its -1

OpenStudy (anonymous):

u in wch class